OpenStudy (gabylovesyou):

Part 1: Create your own quadratic equation that cannot be solved by factoring, but can be solved using the quadratic formula. Identify the values of a, b, and c, and find the solutions using the quadratic formula. Show all work to receive credit. Part 2: Using complete sentences, explain how you know that the equation from Part 1 cannot be solved by factoring, but can be solved by using the quadratic formula.

OpenStudy (anonymous):

Make something in ax^2+bx+c where b^2-4ac is not a neat square root.

OpenStudy (gabylovesyou):

how do i know when its a perfect square ?

OpenStudy (anonymous):

Take several sequential numbers, square them, and then make sure it is between two of those.

OpenStudy (nincompoop):

\[2x^2-3x+\frac{ 1 }{ 2 }=0\]

OpenStudy (gabylovesyou):

3x^2 + 5x + 6 is that good ?

OpenStudy (anonymous):

That doesn't really help unless he knows why. Your polynomial has imaginary roots, Gaby.

OpenStudy (nincompoop):

(x − 5) × (x − 5) (x − 5) × (x − 5) = x2 + -5x + -5x + 25 = x2 + -10x + 25 and x2 + -10x + 25 is a perfect square trinomial

OpenStudy (gabylovesyou):

numbers that i can use ? NO FRACTIONS i havent larned that lol

OpenStudy (nincompoop):

Whenever you multiply a binomial by itself twice, the resulting trinomial is called a perfect square trinomial

OpenStudy (anonymous):

OK, just pick a, b, and c so that b^2-4ac is >0 but not a perfect square number.

OpenStudy (gabylovesyou):

1 2 3 ?

OpenStudy (anonymous):

Not in any order, no.

OpenStudy (nincompoop):

lol good night @Gabylovesyou thanks for taking over, @Numb3r1

OpenStudy (anonymous):

No problem, @nincompoop

OpenStudy (gabylovesyou):

goodnight .-. and @Numb3r1 how about x^2 + x - 4 ?

OpenStudy (anonymous):

That works.

OpenStudy (gabylovesyou):

now to solve it ..

OpenStudy (anonymous):

That works because when you take the quadratic, you have:\[\frac{(b \pm \sqrt{ b^{2}-4ac})}{2a}=\frac{(1 \pm \sqrt{ 1^{2}-(4\times1\times-4)}}{2\times1}=\frac{1 \pm \sqrt{ 17}}{2}\]

OpenStudy (anonymous):

The root 17 would require pretty elaborate factorings which humans don't normally do.

OpenStudy (primeralph):

............I do it.......

OpenStudy (gabylovesyou):

x = 1 +- sqrt(1^2 - 4(1)(-4)) all over 2(1) x = 1 +- sqrt(1 + 16) all over 2 x = 1 +- sqrt(17) all over 2

OpenStudy (gabylovesyou):

is that correct @Numb3r1

OpenStudy (nubeer):

there is usually -b in the starting of the formula not just 'b'

OpenStudy (phi):

As nubeer pointed out the formula is \[x= \frac{(-b \pm \sqrt{ b^{2}-4ac})}{2a} \] notice the "minus b" in front that means your answer should be \[ x= \frac{-1 ± \sqrt{17}}{2}\] for the problem \[ x^2 + x - 4 =0 \] we can test to see if the answer makes the equation true. Do we get 0 ? one of the x answers is -0.5 + sqrt(17)/2= 1.561552813 (1.561552813)^2 + 1.561552813 -4 = a very small number close enough to 0 (we won't get *exactly* zero if we use decimals because the sqrt(17) 's digits go on forever. if we keep x in the form \[ x= \frac{-1 + \sqrt{17}}{2}\] and use algebra, we would get exactly 0.